OpenStudy (luigi0210):

Trig limits help, I know the answer is 1/6, but don't know how they got it.

OpenStudy (luigi0210):

http://prntscr.com/2w2763

OpenStudy (anonymous):

\[\lim_{\theta \rightarrow 0}\frac{ \sin \theta }{ 5 \theta+\tan \theta }\] \[divide~ the~ numerator~ and~ denominator~ by~\theta \] \[=\frac{ \lim_{\theta \rightarrow 0}\frac{ \sin \theta }{\theta } }{ \lim_{\theta \rightarrow0}\left( 5+\frac{ \tan \theta }{\theta } \right) }\] \[=\frac{ 1 }{ 5+1 }=\frac{ 1 }{6 }\]

OpenStudy (luigi0210):

Thank you :D